mass of 100 atom of nitrogen in grams?

Grade:12

5 Answers

Aarti Gupta
8 years ago
We know that mass of 6.023 * 1023 atoms of N2 = Gram atomic mass of N2 = 14g
therefore,
Mass of 1atom of N2 = Gram atomic mass / NA
where,
NA = 6.023 * 1023
Mass of 1atom of N2 = 14/6.023 * 1023
Now,
Mass of 100 atoms of N2 = 14/6.023 * 1023 *100
Mass of 100 atoms of N2 = 2.324 * 10-21g
vindhya
8 years ago
thank you so much
Aarti Gupta
8 years ago
Your welcome dear student.
Prajwal Kavad
8 years ago
mass of one atom of Nis 14/Na, therefore mass of 100 atoms is 1400/N, thats all,,
 
@vindya i would recommend you to read and solve RC muckerjee (mole concept) 1st and 2nd chapters then you will have a strong command over mole concept..
 
thank you
 
vindhya
8 years ago
thanks. if you hav an online link of this book, send me the link.
coz i don’t have that book.
thanks again